# Chatter: A Character Attribution Dataset for Narrative Understanding

### v1

@@ 7 Limitations
We define the character attribution task as a binary classification task, where, given the character-trope pair and the screenplays of the movies in which the character appeared, the model should predict whether or not the character portrayed the trope. This formulation has some limitations. First, the screenplay’s narrative may not exactly resemble the story told by the movie because of tweaks made during the filming process. Publicly available screenplays are rarely the final script but drafts from earlier in the production stage. Second, movies could have visual cues like the non-verbal behavior of the character that are missed by the screenplay text. Lastly, a character can appear in multiple movies, and our dataset does not contain scripts for all of them. These limitations have important implications because the TVTropes contributors and our raters draw their knowledge of the character from all sources. In contrast, the attribution model predicts the attribute labels based only on the screenplay. There could be character-attribute pairs for which the model has insufficient information and could lead to poor recall, as shown by the performance of the baseline models in Table 3.
@@ 3. Choose yes, maybe yes, not sure,
Figure 1: Interface of the annotation task.
Figure 1 shows the annotation interface that the workers use to label character attribution. The annotation task ran from August 11 to September 19, 2024. Throughout the task, we maintained a communication channel for the workers in the TurkerNation Slack Workspace, where we responded to any questions the workers had about the task. We pay \$1.5 for each question which turns out to be \$18/hour because the workers spent an average of 5 minutes per question.
The MTurk workers answer yes, maybe yes, not sure, maybe no or no on each question. We map this ordinal scale to a numeric range by mapping the labels to 2, 1, 0, -1 and -2. Each character-trope pair is annotated by at most three workers. We sum the label values for each sample to get an integer score between -6 and 6. The higher the absolute score, the more agreement between the raters. We drop samples whose absolute integer score was below 3. For the remaining samples, we obtain the annotation confidence $w$ by normalizing the integer score $s$ between 0 and 1: $w=(s-2)/4$. The confidence score takes values 0.25, 0.5, 0.75 and 1. Attribution models can use these scores to weigh their evaluation metrics. Table 3 shows the weighted F1 scores for the baseline models.
@@ Appendix C ScriptsonScreen
Appendix D Prompt
@@
Given below is a movie script enclosed
between the tags <SCRIPT> and </SCRIPT>.
Following that is the definition of the
\$\$TROPE\$\$ trope enclosed between the tags
<TROPE> and </TROPE>.
Read the movie script and based on that
answer yes or no if the character
\$\$CHARACTER\$\$ portrays the \$\$TROPE\$\$
trope.
Also, give a brief explanation as to why
you think the answer is yes or no.
The explanation should follow after the
yes/no answer.
@@
<TROPE>
\$\$DEFINITION\$\$
</TROPE>
We replace \$\$CHARACTER\$\$, \$\$TROPE\$\$, \$\$DEFINITION\$\$ and \$\$SCRIPT\$\$ with the character name, trope name, definition of the trope and the movie script during prompting. We do not ask for an explanation when prompting Llama-3.1-8B-Instruct.
Appendix E Experiments
We used the closed-source Gemini-1.5-Flash and the open-source Llama-3.1-8B-Instruct models to develop the baselines. We used the Gemini API to access Gemini-1.5-Flash. We ran the inference pipeline of Llama-3.1-8B locally on one NVIDIA A40 GPU. The numbers reported in Table 3 were averaged over five runs.

### v2

@@ 8 Limitations
We define the character attribution task as a binary classification task, where, given the character-trope pair and the screenplays of the movies in which the character appeared, the model should predict whether or not the character portrayed the trope. This formulation has some limitations. First, the screenplay’s narrative may not exactly resemble the story told by the movie because of tweaks made during the filming process. Publicly available screenplays are rarely the final script but drafts from earlier in the production stage. Second, movies could have visual cues like the non-verbal behavior of the character that are missed by the screenplay text. Lastly, a character can appear in multiple movies, and our dataset does not contain scripts for all of them. These limitations have important implications because the TVTropes contributors and our raters draw their knowledge of the character from all sources. In contrast, the attribution model predicts the attribute labels based only on the screenplay. There could be character-attribute pairs for which the model has insufficient information and could lead to poor recall, as shown by the performance of the models in Table 5.
@@ 3. Choose yes, maybe yes, not sure,
Figure 1 shows the annotation interface that the workers use to label character attribution.
@@ Appendix C ScriptsonScreen
Appendix D TROPES
There are 445 tropes in total that have at least one character portraying them in our dataset. We could not figure out a systematic way to cluster them because the tropes are very different from each other. Therefore, we simply list the most common tropes and their abbreviated definition. Tables 6 and 7 list the tropes that are portrayed by at least two characters in our dataset.
Trope Definition
1. AdaptationalAttractiveness Plain characters are portrayed as attractive in adaptations
2. AdaptationPersonalityChange Character personality changes during medium adaptations
3. AffablyEvil Charming villain with kindness despite malevolent intentions
4. AntiVillain Heroic goals achieved through questionable or evil means
5. AudienceSurrogate Character audience identifies with for sympathy and relatability
6. BadLiar Character fails to lie convincingly; creates transparent falsehoods
7. BerserkButton Character’s minor trigger causes explosive anger reaction
8. BitchInSheepsClothing Deceptive character appears kind but is secretly villainous
9. BlueIsHeroic Blue signifies heroism through calm, disciplined characters
10. CatchPhrase Repetitive distinctive phrase by a character or category
11. ChildrenAreInnocent Children embody innocence and purity, contrasting adult corruption
12. Cloudcuckoolander Cheerfully eccentric character, detached from reality, unexpectedly wise
13. ControlFreak Obsessively enforces rules, hindering effectiveness and dissent
14. CynicismCatalyst Trauma shifts idealistic character to cynicism and moral ambiguity
15. DarkAndTroubledPast Character’s tragic past shapes their personality and behavior
16. DeadpanSnarker Sarcastic character who critiques and deflates others’ egos
17. Determinator Unyielding persistence towards goals, regardless of challenges
18. DramaQueen Excessively dramatic characters who overreact and seek attention
19. EvenEvilHasStandards Villain shows moral boundaries despite overall remorselessness
20. EveryoneHasStandards Characters uphold personal standards despite their own flaws
21. EvilWearsBlack Villains wear black, symbolizing darkness and aggression
22. FauxAffablyEvil Polite villains hiding true cruelty for manipulation and enjoyment
23. Gaslighting Manipulating perception to induce doubt and confusion
24. GentleGiant Big, kind character defying intimidating appearance; gentle and reliable
25. GoingNative Character embraces new culture, rejects original society
26. GrinOfRage Smiling while angry, used for intimidation or provocation
27. GuileHero Cunning hero uses wit and charm for noble goals
28. HairTriggerTemper Explosive anger at minor irritations; unpredictable and dangerous
29. HeroAntagonist Good antagonist opposing protagonist for noble reasons
30. HiddenDepths Characters unveil unexpected talents, deepening their complexity
31. Hypocrite Authority figures failing to uphold their own ideals
32. ItsAllAboutMe Self-centered character believes world revolves around them
33. Jerkass Self-centered character creating conflict for comedic/dramatic effect
34. JerkassHasAPoint Morally flawed character speaks uncomfortable but true points
35. KarmaHoudiniWarranty Villain faces late justice, satisfying audience’s desire for retribution
36. KickTheDog Character’s cruel act establishes evil, shifts audience sympathy
37. LackOfEmpathy Characters recognize emotions but lack emotional connection
38. LargeHam Flamboyant, over-the-top character adding drama and charisma
39. LaserGuidedKarma Immediate consequences for characters’ actions reinforce moral lessons
40. LivingMacGuffin Person drives quests due to intrinsic value or attributes
41. LoveAtFirstSight Instant deep love between characters upon first meeting
42. LoveInterest Romantic character involved with another, often archetypal roles
43. ManlyTears Stoic male character cries from strong, dignified emotions
44. MeaningfulName Character names reflect traits or roles meaningfully
45. MoralityPet Villain’s bond with innocent character prompts redemption
46. MorphicResonance Characters retain recognizable traits across different forms
47. MyGodWhatHaveIDone Character regrets harmful actions, prompting remorse and conflict
48. Narcissist Character obsessed with self-admiration and validation, often hostile
49. NeverBareheaded Character always wears headgear, never seen bare-headed
50. NiceGirl Kind, friendly character contrasting cynical figures; endearing presence
Table 6: Tropes and their definitions (first part)
Trope Definition
51. NiceGuy Kind, morally good character contrasting with cynicism
52. NiceJobBreakingItHero Hero’s victory causes unintended negative consequences
53. NoCelebritiesWereHarmed Parody characters resembling real celebrities, often with altered names
54. NoSocialSkills Characters lacking social awareness, often blunt but intelligent
55. OhCrap Character realizes impending disaster, leading to panic or horror
56. OnlyOneName Characters known by only one name, no identifiers
57. OnlySaneWoman Rational character amidst chaotic, irrational peers; often frustrated
58. PayEvilUntoEvil Revenge-driven morality blurs hero-villain boundaries
59. PosthumousCharacter Dead character influences plot through memories or narratives
60. ReasonableAuthorityFigure Open-minded leader who evaluates heroes’ claims rationally
61. RebelliousSpirit Character defies norms, follows personal rules, often anti-heroic
62. ShadowArchetype Character reflecting protagonist’s denied flaws, causing conflict and growth
63. SirSwearsALot Character known for excessive swearing, revealing deeper traits
64. StepfordSmiler Cheerful facade hides inner turmoil and psychological issues
65. TheChessmaster Cunning strategist who manipulates events for personal gain
66. TheDon Ruthless crime patriarch with moral codes, protective yet shrewd
67. TheHeart Caretaker and moral compass of the team
68. TheImmune Character immune to disease, pivotal for finding cure
69. TheJerkIndex Characters exhibiting rudeness contrasting with polite behavior
70. TheSociopath Character lacking empathy, manipulative, and morally ambiguous
71. TopHeavyGuy Exaggerated large upper body, skinny legs
72. TragicVillain Sympathetic villain regrets their actions, seeks redemption
73. UncertainDoom Ambiguous fate of a character, survival uncertain
74. UndignifiedDeath Ridiculous, embarrassing deaths blending humor and tragedy
75. UnwittingInstigatorOfDoom Unintentional catalyst for disaster, unaware of their role
Table 7: Tropes and their definitions (second part)
Appendix E Prompt
@@
A character trope is a story-telling
device used by the writer to describe
characters.
Given below is the definition of the
\$TROPE\$ trope enclosed between the
tags <TROPE> and </TROPE>.
Following that is a movie script
enclosed between the
tags <SCRIPT> and </SCRIPT>.
The character "\$CHARACTER\$" appears
in the movie script.
Read the movie script carefully
and based on that answer yes or no
if the character "\$CHARACTER\$"
portrays or is associated with the
\$TROPE\$ trope.
If yes, give a brief explanation.
Answer based only on the movie script.
Do not rely on your prior knowledge.
<TROPE>
\$DEFINITION\$
</TROPE>
@@
Does the character "\$CHARACTER\$"
portray or is associated with
the \$TROPE\$ trope in the above
movie script?
Answer yes or no.
If yes, give a brief explanation.
Do not use MarkDown.
We replace \$CHARACTER\$, \$TROPE\$, \$DEFINITION\$ and \$SCRIPT\$ with the character name, trope name, definition of the trope and the movie script during prompting.